QUESTION 5:

The following information has been extracted from the financial statements of YDI Limited:

Extract of Statement of Comprehensive Income for the year ended 31 December:

                                      2019          2018

                                        $                   $

Sales                       2 000 000 1 600 000

Cost of sales              940 000      800 000

Operating profit       600 000      520 000

Profit before tax      520 000     450 000

Profit after tax           364 000    315 000

Extract of Statement of Financial Position as at 31 December:

Assets                                   2019              2018

                                                    $                     $

Non-current assets          2 000 000 1 400 000

Inventories                            600 000     800 000

Accounts receivable             400 000     400 000

Cash and cash equivalents      2 000          2 000

                                               3 002 000 2 602 000                                                           

                                                        

                                              $                    $

Equity and liabilities

Shareholders’ equity          2 000 000 1 500 000

Long-term loan                      700 000      800 000

Accounts payable                  182 000      142 000

Bank overdraft                       120 000     160 000

                                         3 002 000 2 602 000

Note:

1. All purchases and sales of inventories are on credit.

2. Dividends paid during the year amounted to R218 400.

3. The issued share capital consisted of 500 000 ordinary shares

Required:

5.1 Calculate the following ratios for the year ended December 2019. Where applicable, round off answers to two decimal places.

5.1.1 Operating margin        

5.1.2 Debtors collection period       

5.1.3 Acid test ratio        

5.1.4 Return on equity        

5.1.5 Debt to equity         

5.1.6 Earnings retention ratio        

5.1.7 Earnings per share        

5.2 Suggest two (2) ways in which YDI Limited can improve on its collections from debtors.            

5.3 Comment on the current ratio which dropped from 3.98:1 in 2018 to 3.32:1 in 2019.           

5.4 Recommend two (2) ways in which YDI Limited can improve its profitability.
Answer #1

5.1.1 Operating margin    = operating profit / Net sales

= 600000 / 2000000*100

= 30%

5.1.2 Debtors colllection period = Avererage account recievable / Credit sales *365

* Average account receivable = 400000 + 400000 / 2 = 400000

= 400000 / 2000000 *365 days

= 73 days

5.1.3 Acid test ratio = current assets  / current liabilities

= 600000 + 400000+ 2000 / 182000 + 120000

= 0.33

5.1.4 Return to equity ratio = net income / shareholder's equity

= 450000 / 2000000

= 0.23

5.1.5 Debt to equity ratio = Totat liabilities / shareholders equity

= 1002000 / 2000000   

= 0.50

5.1.6 Earning retention ratio = Net income - dividend paid / net income

= 315000 - 218400 / 315000

    = 0.37

5.2 YDI Limited can improve collection from debtor selling more goods on cash basis rather than credit basis

5.3 current ratio is measures the companies liquidity in YDL Com current ratio decrease from year 2018 to 2019 from 3.98 to 3.32

which indicated that negative which indicate poor position of company.